# MoORE: SVD-based Model MoE-ization for Conflict- and Oblivion-Resistant Multi-Task Adaptation

### v4

@@ 1 Introduction
Parameter-efficient adaptation/fine-tuning [71] plays a central role in the practical deployment of large-scale pre-training models, especially in multi-task learning (MTL) scenarios [80, 61, 11, 62, 42]. Take large language models (LLMs) [82] in natural language processing (NLP) tasks as an example. To increase intrinsic knowledge and maintain generalization power, a pre-trained LLM often needs to learn multiple downstream tasks in different domains simultaneously or sequentially. To achieve multi-task adaptation, parameter-efficient fine-tuning (PEFT) methods like low-rank adaptation (LoRA) [25] and its variants [68, 2, 60, 65, 31] have been proposed. However, the practical applications of these methods are often limited because they often suffer from task conflict and oblivion: $i$) The model adapted for one task often leads to the performance degradation in other tasks (also known as negative transfer [70, 79] or destructive interference [11]). ii)ii) Learning new tasks may result in catastrophic forgetting [62] — the model performance deteriorates severely in previously learned tasks.
@@
(a) (a) An illustration of our model MoE-ization strategy and the corresponding MoORE architecture. (b) The comparison for various multi-task adaptation methods on fine-tuning LLaMA-3.1 8B [20] on the CSR-MTL constructed by nine tasks [9, 8, 45, 5, 54, 77, 53, 59]. MoORE consistently works better than the baselines when the number of tasks is larger than one. (c) Before adaptation, LLaMA-3.1 8B achieves encouraging overall performance (i.e., the gray dashed line) in seven tasks [24, 23, 83, 58, 51, 7, 4, 10]. MoORE mitigates the performance degradation and outperforms the most competitive baseline LoRAMoE [14]. (d) MoORE’s runtime is comparable to that of its competitors. Compared to the original LLaMA-3.1 8B (i.e., the gray dashed line), MoORE increases the inference time moderately.
@@
Essentially, task conflict and oblivion arise because the diversity of different tasks requires the models to adapt their parameters in different directions [72, 74, 13, 26, 15]. Some recent methods [31, 30, 60] combine the Mixture-of-Experts (MoE) architecture [27, 55, 29] with PEFT, mitigating the interferences across different tasks by activating task-specific parameters. Given a layer of a pre-trained model, these methods learn multiple adapters associated with a router in multi-task scenarios. Each adapter may inherit task-specific knowledge, and the router selects/fuses these adapters based on the input data and tasks.
@@
In principle, we call the above strategy “Model MoE-ization” because it converts the original linear layer to an MoE architecture. In this study, we propose a novel model MoE-ization strategy, with the help of the singular value decomposition (SVD), leading to a conflict- and oblivion-resistant multi-task adaptation method. As illustrated in Figure LABEL:fig:scheme, given a weight matrix of a pre-trained model, our method applies SVD [1] to it and introduces a learnable router to adjust its singular values based on tasks and samples. Accordingly, the weight matrix becomes a Mixture of Orthogonal Rank-one Experts (MoORE), where each expert is constructed by the outer product of a left singular vector and the corresponding right one. To further improve the capacity of MoORE, we impose a learnable orthogonal adapter, which is implemented by a Householder reflection adaptation module [75].
@@
Unlike existing methods, which learn some additional low-rank experts without any constraints on their relations, MoORE extracts many rank-one experts intrinsically from the SVD of the original weight matrix. Such a design imposes the orthogonality of the experts, avoiding their information redundancy and undesired interferences. Moreover, similar to existing orthogonal fine-tuning strategies [49, 75], MoORE maintains the column space of the original weight matrix and thus inherit the pre-training ability better. Thanks to the above two properties, MoORE consistently outperforms SOTA methods in various multi-task adaptation scenarios. The results in Figure 4(a) show the superiority of MoORE in mitigating task conflict and oblivion and its competitive inference efficiency.
@@ 2.1 Adapter-based Methods for Multi-Task Adaptation
Multi-task adaptation aims to fine-tune a pre-trained foundation model simultaneously or sequentially in multiple downstream tasks [34, 32]. Focusing on this problem, many adapter-based methods have been proposed [37, 50, 57, 65, 40]. In particular, denote the data of $K$ tasks as $\{\mathcal{D}_{k}\}_{k=1}^{K}$. Given a pre-trained foundation model, whose parameters are denoted as $\bm{\theta}$, multi-task adaptation can often be formulated in the framework of maximum likelihood estimation:
@@
Unlike learning and ensembling different models, adapter-based methods improve the overall model performance by sharing parameters and domain knowledge across various tasks, leading to moderate increases in parameters and complexity. For instance, Hyperformer [44] utilizes a shared hypernetwork to generate task-specific adapters, reducing the number of learnable parameters. Recently, some methods extend LoRA [25] to multi-task adaptation, e.g., MultiLoRA [68], MTL-LoRA [73], HydraLoRA [60], and so on, which learns multiple low-rank adapters to handle diverse tasks.
@@ 2.2 Connections to MoE Architectures
The Mixture-of-Experts (MoE) architecture was initially introduced by the work in [27], which is constructed by multiple specialized networks (called experts) and a router. Given an input data $\bm{x}\in\mathcal{X}$, the MoE derives the output $\bm{y}\in\mathcal{Y}$ as $\bm{y}=\sum_{m=1}^{M}g_{m}(\bm{x})f_{m}(\bm{x})$, where $f_{m}:\mathcal{X}\mapsto\mathcal{Y}$ denotes the $m$-th expert, which achieves a mapping from the sample space $\mathcal{X}$ to the output space $\mathcal{Y}$. $g:\mathcal{\mathcal{X}}\mapsto\mathbb{R}^{M}$ denotes the router, and $g_{m}$ denotes its $m$-th output. The router adjusts the experts’ significance based on input data. When applying a sparse routing strategy, i.e., activating only a few experts for each input [55, 29, 17, 3, 16, 46], the MoE architecture supports building large-scale models while maintaining computational efficiency. Due to its advantages, many large language models, e.g., DeepSeek [12], Grok3, and Qwen3111https://github.com/QwenLM/Qwen3, are built based on MoE, and many efforts have been made to convert well-trained dense models into MoE architectures [28, 22, 52, 47, 78].
As mentioned before, most existing adapter-based multi-task adaptation methods actually “MoE-ize” pre-trained models. Given a weight matrix $\bm{W}$ and its input $\bm{x}$, these methods apply multiple low-rank adapters as experts [76, 14, 35, 41] and add them to the pre-trained models, i.e.,
@@
where $\bm{A}_{m}$ and $\bm{B}_{m}$ are low-rank matrices constructing the $m$-th expert $f_{m}$. For the router $g(\bm{x})$, some attempts have been made to develop advanced routing strategies, e.g., the dynamic routing in AdaMoLE [38] and the token-task hybrid routing in HMoRA [31]. Recently, some new MoE architectures have been proposed, including the asymmetric “Hydra” structure in HydraLoRA [60], the LEMoE [66] for lifelong model editing, and the OMoE [18] for orthogonal output. However, the above methods rely purely on data-driven strategies to determine the experts’ functionality and domain knowledge. Without necessary regularization on the relations across different experts, the experts often suffer from the load imbalance issue, i.e., a limited number of experts are over-trained and applied for multiple tasks, while many experts are seldom used. This issue harms the capacity and generalizability of the models, increasing the risk of task conflict and oblivion.
@@ 3.1 SVD-based Model MoE-ization
  • A hybrid routing strategy: Inspired by HMoRA [31], we adjust the experts’ weights according to input data and tasks, leading to a hybrid routing strategy. Given an input of the $k$-th task, denoted as $\bm{x}^{(k)}\in\mathbb{R}^{d}$, we determines the weight of the $d$-th expert as
@@
  • An orthogonal adapter of input: To further increase the model capacity, we can apply a learnable orthogonal transform to the input, i.e., $\bm{H}\bm{x}$, where the learnable orthogonal transform $\bm{H}$ can be implemented efficiently by the butterfly orthogonal fine-tuning (BOFT) module in [36], the Givens rotation adapter [43], or the Householder reflection adapter (HRA) [75]. In this study, we implement $\bm{R}$ by HRA, which corresponds to the product of $L$ Householder reflections, i.e.,
@@
    whose learnable parameters are $\bm{R}=[\bm{r}_{1},\cdots,\bm{r}_{L}]\in\mathbb{R}^{D\times L}$. Note that applying orthogonal adapters maintains the angles between neurons (i.e., the rows of the weight matrix $\bm{W}$), which helps preserve the knowledge of the pre-trained model when enhancing model capacity [49, 36, 75].
@@
As shown in (7), the overall complexity of MoORE is $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}+L))$. In practice, we set $L,D_{t},D_{s}\ll\min\{D,D_{\text{out}}\}$ to reduce the complexity. Moreover, we can merge the orthogonal adapter into each expert in the inference phase, i.e., obtaining $\bm{V}^{\prime}=\bm{V}^{\top}\bm{H}$, and the complexity further reduces to $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}))$. Figure LABEL:fig:speed shows that the inference efficiency of MoORE is comparable to its competitors.
Table 1: Comparisons for various MoE-based multi-task adaptation methods on their implementations and properties, where $M$ ($D$ in our method) is the number of experts, $k$ is the task index, $\text{GS}(\cdot)$ denotes the Gram-Schmidt orthogonalization [6].
@@
LoRAMoE [14] $\text{Softmax}(\bm{Sx})$ $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$ $M$ $r$ No No
MixLoRA [30] $\text{Top}_{2}(\text{Softmax}(\bm{Sx}))$ $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$ $M$ $r$ No No
MoSLD [81] $\text{Softmax}(\bm{Sx})$ $\{\bm{B}\bm{A}_{m}\}_{m=1}^{M}$ $M$ $r$ No No
HydraLoRA [60] $\text{Softmax}(\bm{Sx})$ $\{\bm{B}_{m}\bm{A}\}_{m=1}^{M}$ $M$ $r$ No No
MTL-LoRA [73] $\text{Softmax}(\bm{\phi}_{k})$ $\{\bm{B}_{m}\bm{\Lambda}_{k}\bm{A}\}_{m=1,k=1}^{M,K}$ MKMK $r$ No No
OMoE [18] $\text{Softmax}(\bm{Sx})$ $\text{GS}(\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M})$ $M$ $r$ Yes No
@@ 3.2 Comparisons with Existing MoE-based Multi-Task Adaptation Methods
  • The design of router. Given an input $\bm{x}\in\mathbb{R}^{D}$, most existing methods leverage a sample-level routing strategy. Typically, they apply a linear projection $\bm{S}\in\mathbb{R}^{M\times D}$ to it and pass the projection result through a softmax operator, leading to a nonnegative and normalized weight vector for $M$ experts. Among them, MixLoRA [30] further applies a sparse routing mechanism — for each input, it only activates the two experts that correspond to the top-2 weights. Instead, MTL-LoRA [73] leverages a task-level routing strategy, which determines the experts’ weights by passing a task-specific embedding (i.e., $\bm{\phi}_{k}\in\mathbb{R}^{M}$, $k\in\{1,\cdots,K\}$ indicates the task index) through a softmax operator. Unlike existing methods, our MoORE considers the task- and sample-level information jointly. The advantage of this hybrid routing strategy is that when the same sample serves for different tasks, MoORE can assign task-specific weights to the experts and thus leverage different domain knowledge accordingly.
  • The design of experts. Most existing methods apply $M$ low-rank adapters as experts, i.e., $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$, where $\bm{B}\in\mathbb{R}^{D_{\text{out}}\times r}$ and $\bm{A}\in\mathbb{R}^{r\times D}$ are two rank-$r$ matrices. To reduce the number of learnable parameters, some methods, e.g., MoSLD [81], HydraLoRA [60], and MTL-LoRA [73], reuse the same $\bm{A}$ or $\bm{B}$ for all experts. MTL-LoRA [73] further introduces a task-specific matrix $\bm{\Lambda}_{k}$ for each expert, where k=1,,Kk=1,...,K indicates the task index. As a result, it creates $MK$ low-rank experts and activates $M$ experts per task. Our MoORE contains $D$ orthogonal rank-one experts $\{\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H}\}_{d=1}^{D}$, in which the orthogonal adapter $\bm{H}$ is shared by all experts. Unlike existing methods, MoORE applies many simple but structured experts. Such a design has several advantages:
@@
      Imposing orthogonality for mitigating task conflict: The experts of MoORE are orthogonal to each other because $\bm{u}_{d}^{\top}\bm{u}_{d^{\prime}}=0$ for all $d\neq d^{\prime}$. The orthogonality ensures that the experts have different functionalities and domain knowledge, without redundant information. In particular, by activating different experts for different tasks, MoORE suppresses the interferences across the tasks in the training phase, thus mitigating the task conflict issue. It is worth noting that O-LoRA [67] also introduces orthogonality constraints within its LoRA adapters. However, there are two notable differences between O-LoRA and MoORE: $i)$ O-LoRA uses a regularization term to pursue orthogonality, which can not strictly make its LoRA adapters orthogonal to each other. ii)ii) O-LoRA does not adopt a MoE architecture. To deal with different tasks, O-LoRA adds new LoRA adapters for each sequentially incoming task, rather than designing and learning a flexible routing mechanism for given and fixed experts. As a result, the model complexity of O-LoRA is linear to the number of tasks. Among existing MoE-based methods, OMoE [18] is the only one imposing orthogonality on experts. However, it applies Gram-Schmidt orthogonalization algorithm [6] to the concatenation of $M$ experts’ output, i.e., $\text{GS}([\bm{B}_{1}\bm{A}_{m}\bm{x},\cdots,\bm{B}_{M}\bm{A}_{m}\bm{x}])$. As a result, imposing orthogonality requires additional $\mathcal{O}(D_{\text{out}}M^{2})$ operations per sample, which is less efficient than ours.
@@
      In single task adaptation scenarios, the work in [75] has shown that maintaining the column space of the weight matrix makes the adapted model inherit the ability of the pre-trained model better, mitigating the oblivion of the previously pre-trained task. In our work, we find that such maintenance is helpful in multi-task adaptation scenarios as well.
  • Computational efficiency. Table 2 shows each method’s learnable parameters and computational complexity. For the methods using the sample-based routing strategy, their routers contain $MD$ learnable parameters. Given a sample, the complexity of the router is $\mathcal{O}(MD)$. For the method [73] using the task-based routing strategy, its router is lightweight, containing $MK$ learnable parameters and determining the weights of its experts with complexity $\mathcal{O}(M)$. For the experts, using $M$ rank-$r$ experts leads to the complexity $\mathcal{O}(M(D_{\text{out}}+D)r)$. Reusing $\bm{A}$ or $\bm{B}$ (e.g., MoSLD [81] and HydraLoRA [60]) and applying sparse routing (e.g., MixLoRA [30]) can reduce the computational complexity significantly. In contrast, introducing additional parameters (e.g., the $\bm{\Lambda}_{k}$ in MTL-LoRA [73]) or operations (e.g., the Gram-Schmidt orthogonalization in OMoE [18]) leads to higher complexity. Existing methods construct an MoE with $M$ rank-$r$ experts, while our MoORE is an MoE with $D$ rank-one experts, whose number of experts is determined by the input dimension and thus much larger than $M$. To improve the computational efficiency of MoORE, we set $D_{s}$ and $D_{t}$ comparable to the rank $r$ and set $L$ comparable to $M$, respectively. As a result, the number of learnable parameters and the complexity of MoORE become comparable to most existing methods.

### v5

@@ 1 Introduction
Parameter-efficient adaptation/fine-tuning xu2023parameter plays a central role in the practical deployment of large-scale pre-training models, especially in multi-task learning (MTL) scenarios zhang2021survey ; vandenhende2021multi ; crawshaw2020multi ; vithayathil2020survey ; lyu2021multi . Take large language models (LLMs) zhao2023survey in natural language processing (NLP) tasks as an example. To increase intrinsic knowledge and maintain generalization power, a pre-trained LLM often needs to learn multiple downstream tasks in different domains simultaneously or sequentially. To achieve multi-task adaptation, parameter-efficient fine-tuning (PEFT) methods like low-rank adaptation (LoRA) hu2021lora and its variants wang2023multilora ; agiza2024mtlora ; tian2024hydralora ; wang2024multimodal ; liaohmora have been proposed. However, the practical applications of these methods are often limited because they often suffer from task conflict and oblivion: $i$) The model adapted for one task often leads to the performance degradation in other tasks (also known as negative transfer wang2019characterizing ; zhang2022survey or destructive interference crawshaw2020multi ). ii)ii) Learning new tasks may result in catastrophic forgetting vithayathil2020survey — the model performance deteriorates severely in previously learned tasks.
(a) An illustration of our SVD-based model MoE-ization strategy and MoORE architecture
(b) Resistance to task conflict
(c) Resistance to task oblivion
(d) Inference efficiency
@@
Figure 1: (a) An illustration of our model MoE-ization strategy and the corresponding MoORE architecture. (b) The comparison for various multi-task adaptation methods on fine-tuning LLaMA-3.1 8B grattafiori2024llama3herdmodels on the CSR-MTL constructed by nine tasks allenai:arc ; clark2019boolq ; OpenBookQA2018 ; Bisk2020 ; sap2019socialiqa ; zellers2019hellaswag ; sakaguchi2021winogrande ; talmor-etal-2019-commonsenseqa . MoORE consistently works better than the baselines when the number of tasks is larger than one. (c) Before adaptation, LLaMA-3.1 8B achieves encouraging overall performance (i.e., the gray dashed line) in seven tasks hendryckstest2021 ; hendrycks2021ethics ; zhou2023instructionfollowingevaluationlargelanguage ; suzgun2022challenging ; rein2024gpqa ; chen2021evaluating ; austin2021program ; cobbe2021gsm8k . MoORE mitigates the performance degradation and outperforms the most competitive baseline LoRAMoE dou2024loramoe . (d) MoORE’s runtime is comparable to that of its competitors. Compared to the original LLaMA-3.1 8B (i.e., the gray dashed line), MoORE increases the inference time moderately.
@@
Essentially, task conflict and oblivion arise because the diversity of different tasks requires the models to adapt their parameters in different directions yadav2023ties ; yu2024language ; deep2024della ; huang2024emr ; du2024parameter . Some recent methods liaohmora ; li2024mixlora ; tian2024hydralora combine the Mixture-of-Experts (MoE) architecture jacobs1991adaptive ; shazeer2017outrageously ; lepikhin2020gshard with PEFT, mitigating the interferences across different tasks by activating task-specific parameters. Given a layer of a pre-trained model, these methods learn multiple adapters associated with a router in multi-task scenarios. Each adapter may inherit task-specific knowledge, and the router selects/fuses these adapters based on the input data and tasks.
@@
In principle, we call the above strategy “Model MoE-ization” because it converts the original linear layer to an MoE architecture. In this study, we propose a novel model MoE-ization strategy, with the help of the singular value decomposition (SVD), leading to a conflict- and oblivion-resistant multi-task adaptation method. As illustrated in Figure 1(a), given a weight matrix of a pre-trained model, our method applies SVD abdi2007singular to it and introduces a learnable router to adjust its singular values based on tasks and samples. Accordingly, the weight matrix becomes a Mixture of Orthogonal Rank-one Experts (MoORE), where each expert is constructed by the outer product of a left singular vector and the corresponding right one. To further improve the capacity of MoORE, we impose a learnable orthogonal adapter, which is implemented by a Householder reflection adaptation module yuan2024bridging .
@@
Unlike existing methods, which learn some additional low-rank experts without any constraints on their relations, MoORE extracts many rank-one experts intrinsically from the SVD of the original weight matrix. Such a design imposes the orthogonality of the experts, avoiding their information redundancy and undesired interferences. Moreover, similar to existing orthogonal fine-tuning strategies qiu2023controlling ; yuan2024bridging , MoORE maintains the column space of the original weight matrix and thus inherit the pre-training ability better. Thanks to the above two properties, MoORE consistently outperforms SOTA methods in various multi-task adaptation scenarios. The results in Figure 1 show the superiority of MoORE in mitigating task conflict and oblivion and its competitive inference efficiency.
@@ 2.1 Adapter-based Methods for Multi-Task Adaptation
Multi-task adaptation aims to fine-tune a pre-trained foundation model simultaneously or sequentially in multiple downstream tasks liu2024mftcoder ; lin2024pemt . Focusing on this problem, many adapter-based methods have been proposed liu2022polyhistor ; raychaudhuri2022controllable ; sung2022vl ; wang2024multimodal ; lu2024prompt . In particular, denote the data of $K$ tasks as $\{\mathcal{D}_{k}\}_{k=1}^{K}$. Given a pre-trained foundation model, whose parameters are denoted as $\bm{\theta}$, multi-task adaptation can often be formulated in the framework of maximum likelihood estimation:
@@
Unlike learning and ensembling different models, adapter-based methods improve the overall model performance by sharing parameters and domain knowledge across various tasks, leading to moderate increases in parameters and complexity. For instance, Hyperformer mahabadi2021parameter utilizes a shared hypernetwork to generate task-specific adapters, reducing the number of learnable parameters. Recently, some methods extend LoRA hu2021lora to multi-task adaptation, e.g., MultiLoRA wang2023multilora , MTL-LoRA yang2025mtl , HydraLoRA tian2024hydralora , and so on, which learns multiple low-rank adapters to handle diverse tasks.
@@ 2.2 Connections to MoE Architectures
The Mixture-of-Experts (MoE) architecture was initially introduced by the work in jacobs1991adaptive , which is constructed by multiple specialized networks (called experts) and a router. Given an input data $\bm{x}\in\mathcal{X}$, the MoE derives the output $\bm{y}\in\mathcal{Y}$ as $\bm{y}=\sum_{m=1}^{M}g_{m}(\bm{x})f_{m}(\bm{x})$, where $f_{m}:\mathcal{X}\mapsto\mathcal{Y}$ denotes the $m$-th expert, which achieves a mapping from the sample space $\mathcal{X}$ to the output space $\mathcal{Y}$. $g:\mathcal{\mathcal{X}}\mapsto\mathbb{R}^{M}$ denotes the router, and $g_{m}$ denotes its $m$-th output. The router adjusts the experts’ significance based on input data. When applying a sparse routing strategy, i.e., activating only a few experts for each input shazeer2017outrageously ; lepikhin2020gshard ; fedus2022switch ; artetxe2021efficient ; du2022glam ; mustafa2022multimodal , the MoE architecture supports building large-scale models while maintaining computational efficiency. Due to its advantages, many large language models, e.g., DeepSeek dai2024deepseekmoe , Grok3, and Qwen3111https://github.com/QwenLM/Qwen3, are built based on MoE, and many efforts have been made to convert well-trained dense models into MoE architectures komatsuzaki2022sparse ; he2024upcycling ; riquelme2021scaling ; nakamura2025drop ; zhang2024bam .
As mentioned before, most existing adapter-based multi-task adaptation methods actually “MoE-ize” pre-trained models. Given a weight matrix $\bm{W}$ and its input $\bm{x}$, these methods apply multiple low-rank adapters as experts zadouri2023pushing ; dou2024loramoe ; liu2024moe ; luo2024moelora and add them to the pre-trained models, i.e.,
@@
where $\bm{A}_{m}$ and $\bm{B}_{m}$ are low-rank matrices constructing the $m$-th expert $f_{m}$. For the router $g(\bm{x})$, some attempts have been made to develop advanced routing strategies, e.g., the dynamic routing in AdaMoLE liu2024adamole and the token-task hybrid routing in HMoRA liaohmora . Recently, some new MoE architectures have been proposed, including the asymmetric “Hydra” structure in HydraLoRA tian2024hydralora , the LEMoE wang2024lemoe for lifelong model editing, and the OMoE feng2025omoe for orthogonal output. However, the above methods rely purely on data-driven strategies to determine the experts’ functionality and domain knowledge. Without necessary regularization on the relations across different experts, the experts often suffer from the load imbalance issue, i.e., a limited number of experts are over-trained and applied for multiple tasks, while many experts are seldom used. This issue harms the capacity and generalizability of the models, increasing the risk of task conflict and oblivion.
@@ 3.1 SVD-based Model MoE-ization
  • A hybrid routing strategy: Inspired by HMoRA liaohmora , we adjust the experts’ weights according to input data and tasks, leading to a hybrid routing strategy. Given an input of the $k$-th task, denoted as $\bm{x}^{(k)}\in\mathbb{R}^{d}$, we determines the weight of the $d$-th expert as
@@
  • An orthogonal adapter of input: To further increase the model capacity, we can apply a learnable orthogonal transform to the input, i.e., $\bm{H}\bm{x}$, where the learnable orthogonal transform $\bm{H}$ can be implemented efficiently by the butterfly orthogonal fine-tuning (BOFT) module in liu2024parameter , the Givens rotation adapter ma2024parameter , or the Householder reflection adapter (HRA) yuan2024bridging . In this study, we implement $\bm{R}$ by HRA, which corresponds to the product of $L$ Householder reflections, i.e.,
@@
    whose learnable parameters are $\bm{R}=[\bm{r}_{1},\cdots,\bm{r}_{L}]\in\mathbb{R}^{D\times L}$. Note that applying orthogonal adapters maintains the angles between neurons (i.e., the rows of the weight matrix $\bm{W}$), which helps preserve the knowledge of the pre-trained model when enhancing model capacity qiu2023controlling ; liu2024parameter ; yuan2024bridging .
@@
As shown in (7), the overall complexity of MoORE is $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}+L))$. In practice, we set $L,D_{t},D_{s}\ll\min\{D,D_{\text{out}}\}$ to reduce the complexity. Moreover, we can merge the orthogonal adapter into each expert in the inference phase, i.e., obtaining $\bm{V}^{\prime}=\bm{V}^{\top}\bm{H}$, and the complexity further reduces to $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}))$. Figure 1(d) shows that the inference efficiency of MoORE is comparable to its competitors.
Table 1: Comparisons for various MoE-based multi-task adaptation methods on their implementations and properties, where $M$ ($D$ in our method) is the number of experts, $k$ is the task index, $\text{GS}(\cdot)$ denotes the Gram-Schmidt orthogonalization bjorck1994numerics .
@@
LoRAMoE dou2024loramoe $\text{Softmax}(\bm{Sx})$ $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$ $M$ $r$ No No
MixLoRA li2024mixlora $\text{Top}_{2}(\text{Softmax}(\bm{Sx}))$ $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$ $M$ $r$ No No
MoSLD zhao2025mosld $\text{Softmax}(\bm{Sx})$ $\{\bm{B}\bm{A}_{m}\}_{m=1}^{M}$ $M$ $r$ No No
HydraLoRA tian2024hydralora $\text{Softmax}(\bm{Sx})$ $\{\bm{B}_{m}\bm{A}\}_{m=1}^{M}$ $M$ $r$ No No
MTL-LoRA yang2025mtl $\text{Softmax}(\bm{\phi}_{k})$ $\{\bm{B}_{m}\bm{\Lambda}_{k}\bm{A}\}_{m=1,k=1}^{M,K}$ MKMK $r$ No No
OMoE feng2025omoe $\text{Softmax}(\bm{Sx})$ $\text{GS}(\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M})$ $M$ $r$ Yes No
@@ 3.2 Comparisons with Existing MoE-based Multi-Task Adaptation Methods
  • The design of router. Given an input $\bm{x}\in\mathbb{R}^{D}$, most existing methods leverage a sample-level routing strategy. Typically, they apply a linear projection $\bm{S}\in\mathbb{R}^{M\times D}$ to it and pass the projection result through a softmax operator, leading to a nonnegative and normalized weight vector for $M$ experts. Among them, MixLoRA li2024mixlora further applies a sparse routing mechanism — for each input, it only activates the two experts that correspond to the top-2 weights. Instead, MTL-LoRA yang2025mtl leverages a task-level routing strategy, which determines the experts’ weights by passing a task-specific embedding (i.e., $\bm{\phi}_{k}\in\mathbb{R}^{M}$, $k\in\{1,\cdots,K\}$ indicates the task index) through a softmax operator. Unlike existing methods, our MoORE considers the task- and sample-level information jointly. The advantage of this hybrid routing strategy is that when the same sample serves for different tasks, MoORE can assign task-specific weights to the experts and thus leverage different domain knowledge accordingly.
  • The design of experts. Most existing methods apply $M$ low-rank adapters as experts, i.e., $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$, where $\bm{B}\in\mathbb{R}^{D_{\text{out}}\times r}$ and $\bm{A}\in\mathbb{R}^{r\times D}$ are two rank-$r$ matrices. To reduce the number of learnable parameters, some methods, e.g., MoSLD zhao2025mosld , HydraLoRA tian2024hydralora , and MTL-LoRA yang2025mtl , reuse the same $\bm{A}$ or $\bm{B}$ for all experts. MTL-LoRA yang2025mtl further introduces a task-specific matrix $\bm{\Lambda}_{k}$ for each expert, where k=1,,Kk=1,...,K indicates the task index. As a result, it creates $MK$ low-rank experts and activates $M$ experts per task. Our MoORE contains $D$ orthogonal rank-one experts $\{\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H}\}_{d=1}^{D}$, in which the orthogonal adapter $\bm{H}$ is shared by all experts. Unlike existing methods, MoORE applies many simple but structured experts. Such a design has several advantages:
@@
      Imposing orthogonality for mitigating task conflict: The experts of MoORE are orthogonal to each other because $\bm{u}_{d}^{\top}\bm{u}_{d^{\prime}}=0$ for all $d\neq d^{\prime}$. The orthogonality ensures that the experts have different functionalities and domain knowledge, without redundant information. In particular, by activating different experts for different tasks, MoORE suppresses the interferences across the tasks in the training phase, thus mitigating the task conflict issue. It is worth noting that O-LoRA wang2023orthogonal also introduces orthogonality constraints within its LoRA adapters. However, there are two notable differences between O-LoRA and MoORE: $i)$ O-LoRA uses a regularization term to pursue orthogonality, which can not strictly make its LoRA adapters orthogonal to each other. ii)ii) O-LoRA does not adopt a MoE architecture. To deal with different tasks, O-LoRA adds new LoRA adapters for each sequentially incoming task, rather than designing and learning a flexible routing mechanism for given and fixed experts. As a result, the model complexity of O-LoRA is linear to the number of tasks. Among existing MoE-based methods, OMoE feng2025omoe is the only one imposing orthogonality on experts. However, it applies Gram-Schmidt orthogonalization algorithm bjorck1994numerics to the concatenation of $M$ experts’ output, i.e., $\text{GS}([\bm{B}_{1}\bm{A}_{m}\bm{x},\cdots,\bm{B}_{M}\bm{A}_{m}\bm{x}])$. As a result, imposing orthogonality requires additional $\mathcal{O}(D_{\text{out}}M^{2})$ operations per sample, which is less efficient than ours.
@@
      In single task adaptation scenarios, the work in yuan2024bridging has shown that maintaining the column space of the weight matrix makes the adapted model inherit the ability of the pre-trained model better, mitigating the oblivion of the previously pre-trained task. In our work, we find that such maintenance is helpful in multi-task adaptation scenarios as well.
  • Computational efficiency. Table 2 shows each method’s learnable parameters and computational complexity. For the methods using the sample-based routing strategy, their routers contain $MD$ learnable parameters. Given a sample, the complexity of the router is $\mathcal{O}(MD)$. For the method yang2025mtl using the task-based routing strategy, its router is lightweight, containing $MK$ learnable parameters and determining the weights of its experts with complexity $\mathcal{O}(M)$. For the experts, using $M$ rank-$r$ experts leads to the complexity $\mathcal{O}(M(D_{\text{out}}+D)r)$. Reusing $\bm{A}$ or $\bm{B}$ (e.g., MoSLD zhao2025mosld and HydraLoRA tian2024hydralora ) and applying sparse routing (e.g., MixLoRA li2024mixlora ) can reduce the computational complexity significantly. In contrast, introducing additional parameters (e.g., the $\bm{\Lambda}_{k}$ in MTL-LoRA yang2025mtl ) or operations (e.g., the Gram-Schmidt orthogonalization in OMoE feng2025omoe ) leads to higher complexity. Existing methods construct an MoE with $M$ rank-$r$ experts, while our MoORE is an MoE with $D$ rank-one experts, whose number of experts is determined by the input dimension and thus much larger than $M$. To improve the computational efficiency of MoORE, we set $D_{s}$ and $D_{t}$ comparable to the rank $r$ and set $L$ comparable to $M$, respectively. As a result, the number of learnable parameters and the complexity of MoORE become comparable to most existing methods.